An ion has a positive charge numerically equal to twice the electronic charge. The mass of the ion is 7000 times that of an electron. The ion is initially stationary and accelerated through a potential difference of 200 V. Calculate the velocity acquired by the ion.
1
Expert's answer
2021-08-04T04:40:48-0400
"\dfrac{1}{2}mv^2=qV,""v=\sqrt{\dfrac{2qV}{m}},""v=\sqrt{\dfrac{2\cdot2\cdot1.6\cdot10^{-19}\ C\cdot200\ V}{7000\cdot9.11\cdot10^{-31}\ kg}}=1.42\cdot10^5\ \dfrac{m}{s}."